OpenStudy (jylarai):

Two numbers are written in scientific notation. The number with a positive exponent is divided by the number with a negative exponent. Describe your result. Explain your answer

OpenStudy (campbell_st):

you need to know about the index law for division \[\frac{x^a}{x^b} = x^{a-b}\] so in terms of scientific notation \[\frac{10^a}{10^{-b}} = 10^{a - (-b)}\] what happens? do you think the power of 10 increases or decreases...?

OpenStudy (jylarai):

Increases

OpenStudy (campbell_st):

yes it does... and why does it increase... that is what you need to explain in your answer...